# Tuning Language Models for Robust Prediction of Diverse User Behaviors

### v1

@@ 3.2 Multitask-Enhanced Anchor Behavior Tuning (A-Tuning)
We divide the instruction fine-tuning data, , into two subsets: , containing labels for anchor behaviors, and , encompassing the rest. In this stage, the LLM is fine-tuned exclusively on . As illustrated in Figure 1, anchor behaviors reflect core patterns of a user’s daily life. Therefore, fine-tuning on anchor behaviors provides the LLM with a foundational understanding of these underlying patterns, enabling accurate next-anchor behavior predictions. Moreover, since the LLM already captures general behavioral knowledge from pretraining, fine-tuning on anchor behaviors facilitates generalization to unseen but semantically related behaviors. Additionally, compared to fine-tuning on the full dataset , this approach mitigates bias toward overrepresented behavior types.
@@
To further improve generalization from anchor to tail behaviors, we proposed a multi-task fine-tuning approach. Beyond learning to predict next behaviors in , we maintain the model’s general task-solving ability, which our experiments later confirmed to be helpful for behavior prediction tasks. To achieve this, we jointly fine-tune the model on an auxiliary dataset , comprising daily user-ChatGPT conversations Zheng et al. (2023). This integrates a conversation generation task alongside the primary behavior prediction task. We control the auxiliary task’s influence by adjusting the size of relative to , using a ratio . Additionally, in practice, we filter out excessively long conversations from to ensure that the prompts for both tasks are comparable in length.
3.3 Difficulty-based Data Selection for Balanced Behavior Tuning (B-Tuning)
@@
In the second stage, we reintroduce the tail behaviors and combine them with to create a class-balanced fine-tuning dataset. Since the LLM fine-tuned during the A-Tuning stage already serves as a good zero-shot predictor for tail behaviors, we believe that a small amount of instruction fine-tuning data covering all behavior types should suffice to build a robust user behavior predictor. However, to achieve this, the quality and informativeness of the selected samples play a crucial role. Specifically, we search for difficult samples from the following two dimensions for effectively fine-tuning the LLM in a few-shot way.
@@
Obtaining Mispredicted Samples. One simple way to measure sample difficulty is by scoring samples with an intermediate model, and those with the wrong predicted labels are more difficult than those correctly predicted Bengio et al. (2009). In our approach, we utilize the LLM fine-tuned during A-Tuning as the scorer and extract mispredicted samples from and for each behavior type.
Contrastive Data Selection. Inspired by contrastive learning Liu et al. (2021), we identify challenging samples by evaluating the distinguishability between their predicted and ground truth labels. Specifically, we select mispredicted samples with lower distinguishability scores , computed as follows:
@@
where denotes the predicted behavior, denotes the ground truth label, and denotes behavior category of , i.e., anchor or tail. For each type of behavior, we only select those mispredicted samples with . This follows the idea that anchor behaviors (tail behaviors) are much more difficult to discriminate from themselves than their counterparts.
@@
Finally, we combine the above two dimensions to measure sample difficulty and select a fixed number of samples for each type of behavior to create the class-balanced fine-tuning dataset§§§Since is usually small, we do not have to score each sample with the LLM to obtain mispredicted samples, which can be time-costly.. By adopting the difficulty-based data selection strategy, we can prompt the LLM to capture more refined behavior knowledge in a sample-efficient manner and significantly enhance its prediction accuracy across tail behaviors without compromising those of anchor behaviors.
@@
4 Experiment
4.1 Experiment Settings
4.1.1 Datasets
We evaluated our model on two real-world user behavior datasets: App usage dataset Li et al. (2022): This open-source dataset records 1,753 users’ interactions across various apps over one week. Grouped by app functions, it includes 24 behaviors reflecting key daily activities. Behavior dataset: This large-scale dataset is sourced from mobile device §§§https://www.honor.com/global/usage logs, capturing user interactions and their corresponding behavioral intentions. After processing, it covers 37 daily behaviors across various scenarios, including learning, work, entertainment, and leisure.
More details about the datasets and preprocessing can be seen in Appendix 3.
4.1.2 Evaluation Protocols and Metrics
To thoroughly evaluate our model, we adopt two evaluation protocols. First, for next behavior prediction Bao et al. (2023); Liao et al. (2024), we split users 8:1:1 into training, validation, and testing sets to assess overall performance. Second, to evaluate tail behaviors, we follow the common practice of long-tailed learning Liu et al. (2019); Shi et al. (2024), construct a balanced test set with 500 samples per behavior, ensuring fair evaluation of tail behaviors. Note this test set is derived from the original Behavior and App datasets, independent of the first protocol. All experiments are repeated five times, with results averaged.
@@ 4.1.2 Evaluation Protocols and Metrics
Overall, we adopt six commonly used metrics. For next behavior prediction, we adopt weighted precision () and weighted recall (), accounting for the overall performance by considering the proportion of behavior categories. Note that the weighted recall () used here corresponds to the common recommendation metric Hit-Rate. For long-tailed learning, following the settings of Liu et al. (2019); Shi et al. (2024), we evaluate four accuracy metrics based on behavior occurrence frequencies: category-average accuracy across all behaviors (), head-category accuracy for behaviors with >5.0% frequency (), medium-category accuracy for those between 1.0% and 5.0% frequency (), and tail-category accuracy for the remaining low-frequency behaviors (). These six metrics assess the model’s robustness across diverse behaviors and its ability to handle imbalanced datasets effectively. Detailed calculation formulas for the six metrics are provided in Appendix B.
4.1.3 Baselines
@@
We selected the following seven representative algorithms to compare with our proposed algorithm, covering traditional methods (SASRec Kang and McAuley (2018) and Bert4Rec Sun et al. (2019)), LLM-enhanced methods (PITuning Gong et al. (2024), LLM-ESR Liu et al. (2024) and AlphaFuse Hu et al. (2025)), LLM-based methods (GPT4o Achiam et al. (2023), A-LLMRec Kim et al. (2024), TALLRec Bao et al. (2023) and LLaRa Liao et al. (2024)). The details of baselines are provided in Appendix C.
4.1.4 Implementation Details
@@
We selected LLama3.1-8B Dubey et al. (2024) as the backbone for our experiments. Our experiments utilized the AdamW optimizer with a cosine annealing learning rate schedule. And we conducted LoRA fine-tuning and parallel training acceleration using the open-source LLM instruction fine-tuning library, llama-factory Zheng et al. (2024). More details about the implementation are in the Appendix D.
App Usage dataset | Behavior Dataset | ||||||||||||
Category | Model | ||||||||||||
Traditional | SASRec | 0.5309 | 0.5759 | 0.2752 | 0.5255 | 0.2567 | 0.1733 | 0.4818 | 0.5507 | 0.1531 | 0.4083 | 0.2159 | 0.0941 |
Bert4Rec | 0.3452 | 0.5400 | 0.0962 | 0.5290 | - | - | 0.1908 | 0.4368 | 0.0290 | 0.2500 | - | - | |
LLM-Enhanced | PITuning | 0.5837 | 0.5133 | 0.2910 | 0.5029 | 0.3721 | 0.1066 | 0.5829 | 0.5057 | 0.2121 | 0.5030 | 0.3545 | 0.0522 |
LLM-ESR | 0.5437 | 0.5906 | 0.2779 | 0.5615 | 0.2467 | 0.1750 | 0.5325 | 0.5750 | 0.1633 | 0.4240 | 0.1772 | 0.0977 | |
AlphaFuse | 0.5621 | 0.6024 | 0.2873 | 0.5706 | 0.2552 | 0.1881 | 0.5428 | 0.5857 | 0.1729 | 0.4335 | 0.1874 | 0.1093 | |
LLM-Based | LLama-NT | 0.5467 | 0.5346 | 0.3736 | 0.5335 | 0.3685 | 0.3000 | 0.5091 | 0.4226 | 0.2433 | 0.4142 | 0.2467 | 0.2040 |
GPT4o | 0.5872 | 0.5678 | 0.4557 | 0.5410 | 0.4642 | 0.4025 | 0.5735 | 0.5660 | 0.3561 | 0.4575 | 0.3703 | 0.3248 | |
A-LLMRec | 0.5908 | 0.6154 | 0.3514 | 0.5815 | 0.3539 | 0.2333 | 0.5545 | 0.5856 | 0.2526 | 0.4567 | 0.3267 | 0.1601 | |
LLaRA | 0.6074 | 0.6256 | 0.4455 | 0.5970 | 0.4468 | 0.3683 | 0.5892 | 0.6099 | 0.3458 | 0.5175 | 0.3661 | 0.2932 | |
TALLRec | 0.6173 | 0.6306 | 0.4397 | 0.6205 | 0.4328 | 0.3580 | 0.5938 | 0.6141 | 0.3425 | 0.5175 | 0.3636 | 0.2908 | |
BehaviorLM | 0.6343 | 0.6364 | 0.5379 | 0.6220 | 0.5240 | 0.5125 | 0.6145 | 0.6261 | 0.4238 | 0.5392 | 0.4353 | 0.3912 | |
Improv. | 2.8% | 0.9% | 18.0% | 0.2% | 12.9% | 27.4% | 3.5% | 2.0% | 19.0% | 4.2% | 17.6% | 20.4% |
4.2 Overall Performance
@@
BehaviorLM consistently outperforms across all metrics on both datasets. Notably, under the long-tailed learning evaluation protocol, BehaviorLM improves Overall Accuracy by up to 18.0% on the App dataset and 19.0% on the Behavior dataset. This performance is driven by its superior ability to predict tail behaviors, surpassing the best baseline by 27.4% (App) and 20.4% (Behavior).
Fine-tuning LLMs for behavior prediction requires addressing the severe long-tailed distribution. Existing LLM fine-tuning methods, such as LLaRA Liao et al. (2024) and TALLRec Bao et al. (2023), which directly fine-tune LLMs on all behavior types, fail to outperform non-tuned GPT4o on Medium and Tail categories. Similarly, hybrid LLM-enhanced methods that integrate LLMs into traditional models lack robustness across both anchor and tail behaviors. BehaviorLM overcomes this challenge through a novel progressive tuning strategy, first mastering anchor behaviors and then generalizing to tail behaviors.
Behavioral knowledge embedded in LLMs significantly enhances behavior prediction. Traditional deep learning-based solutions, such as SASRec Kang and McAuley (2018) and Bert4Rec Sun et al. (2019), struggle to compete with LLM-based prediction models because they cannot leverage this knowledge and perform poorly when fine-tuning data is limited. Additionally, Bert4Rec performs particularly poorly under long-tailed evaluation settings, failing to make accurate predictions on less frequently occurring behaviors (Medium and Tail categories).
4.3 Investigating the Effect of Behavioral Knowledge
Our evaluation in the previous subsection highlights the significant performance improvement brought by the LLM’s behavioral knowledge, as evidenced by the superior performance of LLM-based methods over traditional deep learning approaches. In this subsection, we conduct a more detailed analysis from this perspective. It is well-established that the more parameters an LLM contains, the greater its capacity to store knowledge learned from its pretraining corpus. Therefore, we examine the impact of model size on the LLM’s predictive capability by fine-tuning three versions of BehaviorLM with different backbones: Qwen-1.5B-v2, Llama-8B-v3.1, and Llama-70B-v3.1. Here, the proportion of auxiliary task data is controlled at 5% for all models.
4.3.1 Performance robustness

### v2

@@ 3.2. Multitask-Enhanced Anchor Behavior Tuning (A-Tuning)
We divide the instruction fine-tuning data into two parts: , which contains labels belonging to anchor behaviors, and , which contains the rest. In this stage, we fine-tune the LLM using only . As shown in Figure 1, anchor behaviors represent the core patterns of a user’s daily life. Therefore, after fine-tuning with anchor behaviors, the LLM gains a preliminary understanding of the underlying patterns in user behavior, enabling it to accurately predict the next anchor behavior. Moreover, since the LLM already captures general behavioral knowledge from its pretraining corpus, fine-tuning with anchor behaviors allows it to generalize more easily to other unseen but semantically similar behaviors. This approach, compared to fine-tuning on the entire behavior dataset , helps prevent the LLM from being biased towards a few behavior types that are overrepresented in the data.
@@
To further enhance the LLM’s generalization from anchor to tail behaviors, we propose a multi-task fine-tuning approach. Beyond learning to predict next behaviors in , we maintain the model’s general task-solving capabilities—a strategy our experiments later confirmed to be advantageous for behavior prediction tasks. To achieve this, we simultaneously fine-tune the LLM on another dataset, , derived from daily conversations between users and ChatGPT (Zheng et al., 2023). This effectively integrates an auxiliary task of conversation generation alongside the primary behavior prediction task. We control the impact of the auxiliary task by adjusting the size of , i.e., using a ratio relative to . In practice, we filter out excessively long conversations from to ensure that the prompts from both tasks are comparable in length.
3.3. Difficulty-based Data Selection for Balanced Behavior Tuning (B-Tuning)
@@
In the second stage, we reintroduce the tail behaviors and combine them with to create a class-balanced fine-tuning dataset. Since the LLM fine-tuned during the A-Tuning stage already serves as a good zero-shot predictor for tail behaviors, we believe that a small amount of fine-tuning data covering all behavior types should suffice to build a robust user behavior predictor. However, to achieve this, the quality and informativeness of the selected samples play a crucial role. Specifically, we designed a sample difficulty scoring strategy, searching for difficult samples from the following two dimensions for effectively fine-tuning the LLM in a few-shot way.
@@
Confidence-based difficulty. One simple way to measure sample difficulty is by scoring samples with an intermediate model, and those with the wrong predicted labels are more difficult than those correctly predicted (Bengio et al., 2009). In our implementation, we directly use the fine-tuned LLM from the A-Tuning stage as the sample scorer and compute the prediction confidence for the ground truth behavior . A lower confidence score indicates that the model struggles to model the user’s intention correctly. This forms the first component of our difficulty score: .
Confusion-based penalty. Following the idea of contrastive learning (Liu et al., 2021), we further select difficult samples by considering the distinguishability between their predicted labels and groundtruth labels. We aim to choose those mispredicted samples with a lower distinguishability score (the second component of our difficulty score), which is calculated as follows,
@@
where denotes the predicted behavior, denotes the ground truth label, and denotes behavior category of , i.e., anchor or tail. This follows the idea that anchor behaviors (tail behaviors) are much more difficult to discriminate from themselves than their counterparts.
Finally, we combine the above two dimensions to define the final continuous difficulty coefficient for each sample:
where balances the confidence uncertainty and the semantic confusion penalty, its default value in the experiment was set to 0.5.
Difficulty-Weighted K-Means++ Selection. While the continuous difficulty coefficient provides a nuanced measure of hardness, relying solely on difficulty magnitude for selection creates a new challenge: the selected hard samples might be redundant (e.g., clustering in a specific failure mode). To address this and construct a dataset that is both challenging and representative, we map all samples into a semantic space using the LLM’s embeddings. For each behavior category, we apply the Difficulty-Weighted K-Means++ algorithm to select a fixed number of samples to create the class-balanced fine-tuning dataset§§§Since is usually small, we do not have to score each sample with the LLM to obtain suitable samples, which can be time-costly.. Crucially, we use the normalized difficulty score as the sampling probability weight during cluster initialization. This mechanism ensures that the sampled data covers the diverse semantic landscape of user behaviors while preferentially targeting those instances where the model exhibits high uncertainty or confusion.
Preference Optimization with DPO. Unlike previous A-Tuning Stage that use Supervised Fine-Tuning (SFT), we adopt DPO to further align the model to make fuller use of the selected high-value hard samples. We construct preference pairs where the ground truth is the chosen response , and the incorrect prediction from the A-Tuning model serves as the rejected response . The objective is to minimize the DPO loss:
where is the frozen model from A-Tuning, is the policy model being optimized, and controls the deviation from the reference model.
@@
By adopting the data selection strategy that balances difficulty and diversity, combined with the DPO training paradigm, we can prompt the LLM to capture more refined behavior knowledge in a sample-efficient manner and significantly enhance its prediction accuracy across tail behaviors without compromising those of anchor behaviors.
@@
4. Experiment
4.1. Experiment Settings
4.1.1. Datasets.
We evaluated our model on two real-world user behavior datasets:
Honor behavior dataset: This large-scale dataset is derived from mobile device §§§Corresponding Author.usage logs. When users interact with their mobile phones, various types of logs are generated, desensitized, and reported with user consent. We select 37 daily behaviors that are reliably extracted from raw logs and also cover broad life scenarios, including activities related to learning, work, entertainment, leisure, etc. The dataset spans from March 1, 2024, to April 29, 2024, and consists of over 50 million behavior events from 24,133 anonymous users. We preprocess the dataset and construct samples in the format of “(last-20 events, next event)”. Since our target is fine-tuning the LLM instead of training from scratch, we only randomly select a subset (200,000) of data for experiments.
App usage dataset (Li et al., 2022): This dataset is an open-source resource that captures 1753 user interactions across various apps within one week. Given the large number of apps and the overlap in functionality among many of them, we processed the dataset further by merging apps with similar purposes. For example, Douyin and Kuaishou were grouped into a ”watching videos” category. After similar preprocessing as Honor dataset, this dataset consists of 71 behavior events belonging to 24 categories.
The characteristics of both datasets are presented in Table 2, and the detailed information, including behavior types for the Honor behavior dataset, is provided in Appendix A.
| Dataset | # User | # Behavior Type | # Sample |
| Honor Behavior Dataset | 24,133 | 37 | 200,000 |
| App Usage Dataset | 1,753 | 24 | 50,000 |
4.1.2. Evaluation Protocols and Metrics
To comprehensively evaluate our model, we consider two different evaluation protocols. First, we follow the common practice of next behavior prediction (Bao et al., 2023; Liao et al., 2024) to evaluate the overall prediction performance of the model. First, we divided the users in an 8:1:1 ratio and applied consistent data processing methods to generate the training set, validation set, and test set samples. Second, since we aim to evaluate whether the model can robustly predict across both anchor and tail behaviors, we also follow common practice of long-tailed learning (Liu et al., 2019; Shi et al., 2024) to construct an additional test set, where each behavior category contains equally 500 samples, ensuring that rarely-occurred tail behaviors are sufficiently evaluated compared with the first next behavior prediction protocol that reflects the real-world distribution of behavior categories. Note that this additional test set is extracted from the original Hornor dataset and App dataset, which were not used in the first protocol. For each evaluation protocol, we repeat experiments five times and report mean values.
@@ 4.1.2. Evaluation Protocols and Metrics
Overall, we adopt six commonly used metrics. The next behavior prediction evaluation considers weighted precision () and weighted recall (), accounting for the overall performance by considering the proportion of behavior categories. Note that the recall calculation () used here is equivalent to the widely adopted recommendation metric Hit-Rate. The long-tailed learning evaluation considers four specific accuracy metrics according to the occurrence frequency of different behaviors. These are category-average accuracy for all behavior types (denoted as ), head-category accuracy for whose frequency larger than 5.0% (denoted as ), medium-category accuracy for whose frequency larger than 1.0% and lower than 5.0% (denoted as ), and tail-category accuracy for the rest behavior types (denoted as ), following the settings of (Liu et al., 2019; Shi et al., 2024). By utilizing these metrics, we focus on whether the model can make robust predictions across diverse behaviors and better assess its predictive performance on imbalanced behavioral datasets. The specific calculation formula for the above six metrics is listed in Appendix B.
4.1.3. Baselines
@@
We selected the following seven representative algorithms to compare with our proposed algorithm, covering traditional methods (SASRec (Kang and McAuley, 2018) and Bert4Rec (Sun et al., 2019)), LLM-enhanced methods (PITuning (Gong et al., 2024), LLM-ESR (Liu et al., 2024) and AlphaFuse (Hu et al., 2025)), LLM-based methods (GPT4o (Achiam et al., 2023), A-LLMRec (Kim et al., 2024), TALLRec (Bao et al., 2023), LLaRa (Liao et al., 2024)) and CoLLM (Zhang et al., 2025b). The details of baselines are provided in Appendix C.
4.1.4. Implementation Details.
@@
We selected LLama3.1-8B (Dubey et al., 2024) as the backbone for our experiments. To ensure flexibility in model testing, we designed three distinct instruction formats, which are randomly sampled during both training and testing. Our experiments utilized the AdamW optimizer with a cosine annealing learning rate schedule, setting the warm-up proportion to 0.1. The maximum learning rate for cosine annealing was set to 1e-4, while both the minimum and initial warm-up learning rates were set to 1e-6. We conducted LoRA fine-tuning and parallel training acceleration using the open-source LLM instruction fine-tuning library, llama-factory (Zheng et al., 2024). All experiments were performed with a maximum of 8 training epochs and a batch size of 8, selecting the best-performing model on the validation set for testing. Detailed formats of the three instruction types are provided in the Appendix D.
| (a) App Usage Dataset | |||||||
| Category | Model | ||||||
| Traditional | SASRec | 0.5309 | 0.5759 | 0.2752 | 0.5255 | 0.2567 | 0.1733 |
| Bert4Rec | 0.3452 | 0.5400 | 0.0962 | 0.5290 | - | - | |
| LLM-Enhanced | PITuning | 0.5837 | 0.5133 | 0.2910 | 0.5029 | 0.3721 | 0.1066 |
| LLM-ESR | 0.5437 | 0.5906 | 0.2779 | 0.5615 | 0.2467 | 0.1750 | |
| AlphaFuse | 0.5621 | 0.6024 | 0.2873 | 0.5706 | 0.2552 | 0.1881 | |
| LLM-Based | LLama-NT | 0.5467 | 0.5346 | 0.3736 | 0.5335 | 0.3685 | 0.3000 |
| GPT4o | 0.5872 | 0.5678 | 0.4557 | 0.5410 | 0.4642 | 0.4025 | |
| A-LLMRec | 0.5908 | 0.6154 | 0.3514 | 0.5815 | 0.3539 | 0.2333 | |
| LLaRA | 0.6074 | 0.6256 | 0.4455 | 0.5970 | 0.4468 | 0.3683 | |
| CoLLM | 0.6053 | 0.6264 | 0.4478 | 0.6032 | 0.4423 | 0.3726 | |
| TALLRec | 0.6173 | 0.6306 | 0.4397 | 0.6205 | 0.4328 | 0.3580 | |
| BehaviorLM | 0.6440 | 0.6385 | 0.5514 | 0.6375 | 0.5308 | 0.5265 | |
| Improv. | 4.3% | 1.3% | 21.0% | 2.7% | 14.3% | 30.8% | |
| (b) Behavior Dataset | |||||||
| Category | Model | ||||||
| Traditional | SASRec | 0.4818 | 0.5507 | 0.1531 | 0.4083 | 0.2159 | 0.0941 |
| Bert4Rec | 0.1908 | 0.4368 | 0.0290 | 0.2500 | - | - | |
| LLM-Enhanced | PITuning | 0.5829 | 0.5057 | 0.2121 | 0.5030 | 0.3545 | 0.0522 |
| LLM-ESR | 0.5325 | 0.5750 | 0.1633 | 0.4240 | 0.1772 | 0.0977 | |
| AlphaFuse | 0.5428 | 0.5857 | 0.1729 | 0.4335 | 0.1874 | 0.1093 | |
| LLM-Based | LLama-NT | 0.5091 | 0.4226 | 0.2433 | 0.4142 | 0.2467 | 0.2040 |
| GPT4o | 0.5735 | 0.5660 | 0.3561 | 0.4575 | 0.3703 | 0.3248 | |
| A-LLMRec | 0.5545 | 0.5856 | 0.2526 | 0.4567 | 0.3267 | 0.1601 | |
| LLaRA | 0.5892 | 0.6099 | 0.3458 | 0.5175 | 0.3661 | 0.2932 | |
| CoLLM | 0.5928 | 0.6125 | 0.3493 | 0.5213 | 0.3674 | 0.2964 | |
| TALLRec | 0.5938 | 0.6141 | 0.3425 | 0.5175 | 0.3636 | 0.2908 | |
| BehaviorLM | 0.6317 | 0.6402 | 0.4389 | 0.5558 | 0.4506 | 0.3978 | |
| Improv. | 6.4% | 4.3% | 23.3% | 6.6% | 21.7% | 22.5% | |
4.2. Overall Performance
@@
BehaviorLM consistently achieves the best performance on both datasets, as evaluated by various metrics. Notably, under the long-tailed learning evaluation protocol, BehaviorLM shows an improvement of up to 21.0% in Overall Accuracy for the App dataset and 23.3% for the Behavior dataset. This is primarily driven by a significant improvement in predicting tail behaviors, where BehaviorLM outperforms the best baseline by 30.8% (App) and 22.5% (Behavior).
Fine-tuning LLMs for behavior prediction requires addressing the severe long-tailed distribution found in empirical data. Existing LLM fine-tuning approaches for behavior prediction struggle with modeling this long-tailed distribution. LLM-based models such as LLaRA (Liao et al., 2024) and TALLRec (Bao et al., 2023), which directly fine-tune the LLM on behavior data across all types, fail to outperform non-tuned GPT4o on the Medium and Tail categories. Similarly, LLM-enhanced methods that incorporate LLM-encoded knowledge into traditional models still do not achieve robust learning across both anchor and tail behaviors. In contrast, BehaviorLM employs a novel progressive tuning approach, enabling the LLM to first master predicting anchor behaviors and then generalize to the remaining tail behaviors.
The behavioral knowledge stored in LLMs proves highly beneficial for predicting user behaviors. Traditional deep learning-based solutions, such as SASRec (Kang and McAuley, 2018) and Bert4Rec (Sun et al., 2019), struggle to compete with LLM-based prediction models because they cannot leverage this knowledge and perform poorly when fine-tuning data is limited. Additionally, Bert4Rec performs particularly poorly under long-tailed evaluation settings, failing to make accurate predictions on less frequently occurring behaviors (Medium and Tail categories).
4.3. Investigating the Effect of Behavioral Knowledge
Our evaluation in the previous subsection highlights the significant performance improvement brought by the LLM’s behavioral knowledge, as evidenced by the superior performance of LLM-based methods over traditional deep learning (DL) approaches. In this subsection, we conduct a more detailed analysis from this perspective. It is well-established that the more parameters an LLM contains, the greater its capacity to store knowledge learned from its pretraining corpus. Therefore, we examine the impact of model size on the LLM’s predictive capability by fine-tuning three versions of BehaviorLM with different backbones: Qwen-1.5B-v2, Llama-8B-v3.1, and Llama-70B-v3.1. Here, the proportion of auxiliary task data is controlled at 5% for all models.
4.3.1. Performance robustness.